For this problem do I sub each given in the place of f or for x?
Evaluate each function for the value specified.
f (x) = x2 - 7x + 10; find (a) f (0), (b) f (5), and (c) f (-2).

Answer by tutorcecilia(2152)

f (x) = x2 - 7x + 10 [yes, just substitute and solve the equation]
.
f (0) = (0)2 - 7(0) + 10
f(0)=10
.
f (5) = (5)2 - 7(5) + 10
f (5) = 25-35 + 10
f(5)= -10+10
f(5)= 0
.
f (-2) = (-2)2 - 7(-2) + 10
f(-2)=4+14+10
f(-2)=28